# UniMLVG: Unified Framework for Multi-view Long Video Generation with Comprehensive Control Capabilities for Autonomous Driving

### v2

@@ Abstract
The creation of diverse and realistic driving scenarios has become essential to enhance perception and planning capabilities of the autonomous driving system. However, generating long-duration, surround-view consistent driving videos remains a significant challenge. To address this, we present UniMLVG, a unified framework designed to generate extended street multi-perspective videoscise control. By integrating single- and multi-view driving videos into the training data, our approach updates cross-frame and cross-view modules across three stages with different training objectives, substantially boosting the diversity and quality of generated visual content. Additionally, we employ the explicit viewpoint modeling in multi-view video generation to effectively improve motion transition consistency. Capable of handling various input reference formats (e.g., text, images, or video), our UniMLVG generates high-quality multi-view videos according to the corresponding condition constraints such as 3D bounding boxes or frame-level text descriptions. Compared to the best models with similar capabilities, our framework achieves improvements of 21.4% in FID and 36.5% in FVD.
@@ 1 Introduction
From these frontier explorations, the main demands of effective generated driving videos include long-term multi-view consistency, condition-based controllability, and diversity. However, existing generative algorithms fail to satisfy these requirements simultaneously. For instance, DriveGAN [24] and DriveDreamer [49] introduce action-based autoregressive techniques to generate next-frame driving images. But, they both overlook the importance of multi-view perspectives in autonomous driving, limiting their ability to generate comprehensive multi-view data. MagicDrive [12] addresses this by incorporating 3D control information and cross-view modules to create multi-view images. Although they leverage the approach [55] to produce videos, their outputs are short and lack temporal consistency. Recent efforts [50, 30, 23] focus on generating controllable, multi-view videos over longer sequences. However, due to the limitations of single-objective training and small-scale datasets, they struggle to achieve temporal consistency and diversity. Importantly, the aforementioned works depend on scene-level descriptions to generate corresponding videos without fine-grained text conditions.
To achieve further breakthroughs, this paper proposes a novel generative framework UniMLVG to obtain highly consistent and controllable multi-view videos. Built upon a DiT-based image generation model [10], UniMLVG incorporates temporal and cross-view modules to capture dynamic sequences and information across multiple viewpoints. To enhance consistency, the multi-task training objective constraints and explicit perspective modeling are considered in our method. For diversity of generated instances, over 1,000 hours driving scene data with single-view and multi-view formats is utilized for model training. In terms of multi-condition control, we integrate both 3D conditions—such as 3D bounding boxes and high-definition maps (HDmaps)—and image-wise textual descriptions. In addition, our UniMLVG effectively supports text-to-video (T2V), image-to-video (I2V), and video-to-video (V2V) generation by using a progressive training strategy and demonstrates exceptional performance in both text-based and 3D condition-based editing, as shown in Fig. 1. In summary, the primary contributions involve:
@@ 2.2 Multi-view Video Generation
Compared to standard video generation, multi-view video generation remains relatively underexplored, due to the challenges in ensuring the consistency across perspectives and time series initial trials simply adopts the video diffusion architecture. For instance, MagicDrive [12] encodes high-level controls such as text and bounding boxes independently and flatten those conditions into an embedding sequence. In contrast, DreamForge [30] employs a trainable ControlNet [61] to fuse multi-modality condition. Drive-WM [50] generates videos from a few views using a diffusion model in the first stage and then utilizes multi-view factorization to predict additional views conditioned on the pre-generated ones. DiVE [23] leverages the DiT structure within its diffusion model and applies a view-inflated attention mechanism to compute attention across features from all viewpoints. Extending the design of [49], DriveDreamer-2 [62] concatenates all images per frame into a large single image, and only cross-frame modules are adopted in the forward pass. Despite these approaches, they usually rely on a reference frame to generate subsequent content and the quality of the videos significantly degrades as the sequence length increases. Additionally, they struggle to address inconsistencies across different views and timestamps. Based on the aforementioned issues, we introduce our unified model, which can provide precise control, generate high-quality videos and guarantee the spatial and temporal consistency.
@@ 3 Method
Figure 2: Overall framework of the model. left: The encoded reference frames are concatenated with the noise latent as video latent and fed into $N$ UniMLVG blocks with the text latent. The diverse conditions including image-level descriptions, camera pose, and 3D conditions are injected into each UniMLGV block and interact with the video latent to guide the generated contents. Finally, the model outputs the subsequent frames, which can then be used as the reference frame for the next autoregressive generation. Note that our model can produce driving video based on those conditions only, where the reference frames are not required. right: Details of the UniMLVG block. A UniMLVG block comprises three distinct sub-blocks to perform attention across different dimensions, while the different conditions are integrated into the video latent in different positions during the forward passing.
Fig. 2 illustrates the overall architecture of our proposed UniMLVG. Our method mainly enhances the DiT-based image generation model [10] with two additional modules: the temporal module $\mathcal{T}$ and the cross-view module ${}\mathcal{C}$. UniMLVG is capable of generating extended street-view videos and performing text-based and 3D-conditioned editing by using a multi-task, multi-condition, and multi-stage training strategy across diverse datasets. And, the integration of different street-view datasets (single/multi-view) with the corresponding modeled viewpoints is utilized to achieve the generation of consistent, long-duration multi-view videos.
@@ 3.1 Unified Framework
We choose the DiT-based model [10] as the backbone for image generation over alternatives like [38, 3, 11, 32] and extend the MM-DiT block, shown in the rigth of Figure 2 to our UniMLVG block to achieve an optimal balance of generation quality, scalability, model size, and flexibility in text-based control. For our diffusion model’s optimization, we employ the conditional rectified flows [26, 27, 10] loss as the primary objective.
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|v_{\theta}(z_{t},t,c)-(z_{0}-\epsilon)\right\|^{2}\right],$
where $v$ is the model, $\theta$ is the parameters, $t$ is the timestep and $c$ is the condition.
@@ 3.2 Multi-task
We observe that when using only the video prediction task, the quality of the generated video deteriorates significantly after a few autoregressive iterations. We believe this is due to the model’s excessive reliance on reference frames, which leads to accumulated autoregressive errors. To address this issue, we propose a multi-task training strategy to improve long-term video quality and coherence. Specifically, we design four training tasks: video prediction (VP), image prediction (IP), video generation (VG), and image generation (IG). In VP, we utilize the embeddings of the first $k$ frames from $n$ viewpoints as reference frames, setting their timestep to 00. In IP, we mask 50% of the reference frames, requiring the model to use the remaining discrete reference information to generate images at the masked positions. In VG, the model generates the next $l$ frames of a multi-view video based solely on the given conditions, without any reference frames. In IG, we drop the temporal module to prevent the model from overly relying on temporal continuity, thus preserving its ability to maintain cross-view consistency. During training, these tasks are executed by sampling a mask $M\in\mathbb{R}^{T\times V}$ based on a predefined ratio, with the loss calculation excluding predictions of reference frames, as described below:
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|(1-M)\odot(v_{\theta}(z_{t},t,c)-(z_{0}-\epsilon))\right\|^{2}\right].$
Additionally, our multi-task approach differs from that of DriveDreamer-2 [62] in several aspects. First, UniMLVG utilizes multiple reference frames to achieve more diverse combinations. Second, unlike DriveDreamer-2, which is limited to a forward-facing reference frame, UniMLVG randomly selects viewpoints in IP. Furthermore, we incorporate the IG task to enhance performance.
@@ 3.3 Multi-condition
To help the model grasp the physical dynamics of autonomous driving scenes, we introduce both local conditions (such as 3Dboxes and HDmaps) and global conditions (including view-specific text descriptions). Notably, we are the first to explicitly model camera parameters to incorporate perspective condition for this task.
Local Conditions. We unify local conditions as image-based conditions. Similar to the approach used in T2I-Adapter [31], we introduce a lightweight image adapter for efficient processing. By projecting 3Dboxes or HDmaps and mapping the instance identity to the color space, we generate sparse images $I_{l}\in\mathbb{R}^{T\times V\times H^{\prime}\times W^{\prime}\times 3}$, matching the original image size. These conditional images are concatenated and input into the adapter to obtain multi-level features $C_{l}\in\mathbb{R}^{k\times T\times V\times H\times W\times C}$, which are then added to the latent representation at the corresponding layer to integrate the conditional information.
@@
Figure 3: Field of view overlap between cameras over time. It can reduce the uncertainty in generating new scene frames.
Perspective Modeling. Apparently, there exist overlaps in the image contents between the adjacent views at the same moment and between the same view at different time. Moreover, as shown in Figure 3, there is also a significant field of view overlap between the current viewpoint of the moving vehicle and one of previous viewpoint. All of these overlaps help to reduce the uncertainty in generating new scene images. Meanwhile, modeling camera poses explicitly has proven effective in recent works [13, 40] that generate multi-view images of static scenes. We believe this is also applicable for generating consistent dynamic driving scenes. Specifically, we set the optical center of the forward-facing camera in the first frame as the origin to establish a unified coordinate system for all viewpoints. The camera rays are encoded to produce ray features and subsequently added to the image latent representation. The formula for ray coordinate encoding is as follows, where $j$ represents the encoding dimension (set to $8$ in this paper), and $p$ represents the coordinate value.
$Enc(v)=$ $MLP(\sin(2^{0}\pi p),\cos(2^{0}\pi p),\sin(2^{1}\pi p),$
$\cos(2^{1}\pi p),\dots,\sin(2^{j-1}\pi p),\cos(2^{j-1}\pi p)).$
@@
Ours $\surd$ $\surd$ 20s 8.8 60.1 22.50 70.81 29.12
@@ 3.4 Multi-stage
Our goal is to utilize a wide range of currently available datasets, including both single-view and multi-view street scene videos, to impart driving scene priors to the model. Considering that single-view and multi-view data differ in the number of viewpoints and annotation information, we implement a multi-stage training strategy to ensure stable and efficient model convergence.
@@
Stage II. Infusing the model with the ability to generate from multiple viewpoints and effectively follow conditional inputs. We train $\mathcal{C}$ using several multi-view datasets [5, 43, 52]. These datasets provide multi-view videos, camera calibration parameters, and 3D annotations, allowing the model to generate diverse multi-view videos. During this phase, only the temporal and cross-view modules are trained, while the backbone remains frozen.
Stage III. To further improve generation quality, we perform full fine-tuning in this phase. With the model’s generation capabilities well-developed from the first two stages, this phase typically requires a minimal number of training epochs. Therefore, this stage is optional.
@@ 4.1 Experiment Details
Datasets. We use the single-view dataset OpenDV-Youtube [57] and the multi-view datasets nuScenes [5], Waymo [43], and Argoverse2 [52]. OpenDV-Youtube is exclusively used for pretraining, while nuScenes, Waymo, and Argoverse2 are divided into training and validation sets following their original splits. The training duration totals $1,498$ hours, comprising $1,486$ hours from OpenDV-Youtube, $4.6$ hours from nuScenes, $4.4$ hours from Waymo, and $3.1$ hours from Argoverse2. We leverage available dataset annotations, including 3Dboxes, HDmaps and camera parameters. Meanwhile, nuScenes with $12$ Hz interpolated annotations [2] is used. Additionally, text descriptions for all frames and views are generated at $2$ Hz using [48].
Evaluation Metrics. To assess the effectiveness of our method in terms of realism, continuity, and precise control, we selected four key metrics to compare against existing multi-view image and video generation methods. For realism, we use the widely recognized Fréchet Inception Distance (FID) [17]. To estimate temporal coherence in our videos, we measure consistency using Fréchet Video Distance (FVD) [45]. For a fair comparison, we conduct evaluation following [12], using $150$ scenes from the nuScenes validation set, with $6$ viewpoints per scene and $16$ frames per viewpoint, totaling $900$ videos for FVD computation. Additionally, we sample $10,000$ images from these $900\times 16$ frames to calculate FID. For controllability, we evaluate two perception tasks: 3D object detection [28] and BEV segmentation [64], following the approach MagicDrive [12].
@@ 4.2 Experiment Results
Quantitative Results. We report quantitative experimental metrics on the nuScenes validation set, as shown in Tab. 1. Overall, our model achieves quite promising results, with a significant improvement of 21.4% in FID compared to the second-best method. Firstly, this is attributed to the powerful text-to-image model [10] which is used as our backbone. Additionally, as demonstrated in the ablation study 4.3, the multi-task and multi-stage training strategies make significant contributions. In terms of temporal consistency, we achieve a 36.5% improvement over DiVE [23], which also uses the DiT [18] architecture, and obtain the second place in FVD, slightly lower than DriveDreamer-2 [62]. However, DriveDreamer-2 can only produce 7s video while UniMLVG can generate longer video up to $20$s, almost three times as longer as DriveDreamer-2. Additionally, unlike DriveDreamer-2 [62], UniMLVG does not stitch multiple views into a single image for generation, which reduces memory usage and training time. On contrast, rur cross-view module can easily handle the different numbers of viewpoints presented in different datasets. In terms of condition adherence, UniMLVG achieve the SOTA results, improving $\text{mIoU}_{road}$ and $\text{mIoU}_{vehicle}$ by 8.8% and 7.8%, respectively, over the second-best method. For mAP, we outperform all other methods except DiVE [23]. In contrast to DiVE, UniMLVG does not assign distinct classifier-free guidance scales for each condition, significantly reducing inference time. Furthermore, rather than using the parameter-heavy ControlNet [61], UniMLVG employs a lightweight adapter, enabling more flexible and efficient condition integration.
@@
$\surd$ 79.30 12.80
$\surd$ 79.18 11.86
$\surd$ $\surd$ 82.13 11.91
$\surd$ $\surd$ $\surd$ 82.76 11.74
$\surd$ $\surd$ $\surd$ 76.52 11.71
@@
Controllability. Our model supports 3D condition control as well as text-based control capabilities. In Figure 1(c), we generate realistic scenes using 3D conditions obtained from the simulation engine CALAR [9]. Notably, different from style transfer methods, our generation is not strictly constrained to transforming the content of simulated scenes. Instead, our model takes the general 3D information such as 3D bounding boxes and HDMaps as inputs and leverages real-world knowledge to generate plausible scenes that align with the distribution of the training data. Specifically, elements such as roads, mountains, and trees are generated with appearances that more accurately reflect how they would look in the real world. This task is highly meaningful because, in simulation, we can create various corner cases, and the model can reimagine these scenarios as real-world conditions, thereby addressing the scarcity of such data. For textual control, the model illustrates strong zero-shot generalizability and impressive editing capabilities under extreme transformations. As shown in Figures 1(d) and 4, even though the nuScenes dataset does not contain any snowy, our model can still edit the weather of nuScene multi-view videos in both daytime and nightime. Additionally, the generated videoes accurately present real-world details such as snow accumulating on the roadside, bare trees during winter, and reflective road surfaces due to rain.

### v3

@@ Abstract
The creation of diverse and realistic driving scenarios has become essential to enhance perception and planning capabilities of the autonomous driving system. However, generating long-duration, surround-view consistent driving videos remains a significant challenge. To address this, we present UniMLVG, a unified framework designed to generate extended street multi-perspective videoscise control. By integrating single- and multi-view driving videos into the training data, our approach updates a DiT-based diffusion model equipped with cross-frame and cross-view modules across three stages with multi training objectives, substantially boosting the diversity and quality of generated visual content. Importantly, we propose an innovative explicit viewpoint modeling approach for multi-view video generation to effectively improve motion transition consistency. Capable of handling various input reference formats (e.g., text, images, or video), our UniMLVG generates high-quality multi-view videos according to the corresponding condition constraints such as 3D bounding boxes or frame-level text descriptions.ompared to the best models with similar capabilities, our framework achieves improvements of 48.2% in FID and 35.2% in FVD.
@@ 1 Introduction
Recent advancements in generating driving videos highlight several key requirements: long-term multi-view consistency, condition-based controllability, and diversity. However, existing generative algorithms fail to satisfy these requirements simultaneously. For instance, DriveGAN [24] and DriveDreamer [50] introduce action-based autoregressive techniques to generate next-frame driving images. Nevertheless, they both overlook the importance of multi-view perspectives in autonomous driving, limiting their ability to generate comprehensive multi-view data. MagicDrive [12] addresses this by incorporating 3D control information and cross-view modules to create multi-view images. Although they leverage the approach [56] to produce videos, the outputs are short in duration and lack temporal consistency. More recent efforts [51, 30, 23] have shifted focus toward generating controllable, multi-view videos over longer sequences. However, due to the constraints of single-objective training and small-scale datasets, these methods still struggle with achieving both temporal consistency and diversity. Notably, all of these works rely on scene-level descriptions to generate corresponding videos, but they neglect the inclusion of fine-grained textual conditions, which limits their overall quality and controllability.
To address the limitations outlined above, this paper introduces a novel generative framework, UniMLVG, designed to generate highly consistent and controllable multi-view long videos. Building on a DiT-based image generation model [10], UniMLVG integrates temporal and cross-view modules to capture dynamic sequences and multi-view information. To mitigate error accumulation caused by long-term autoregressive generation, we propose a multi-task training objective. Additionally, to ensure consistent motion across views, we introduce explicit perspective modeling. Unlike previous methods that directly encode camera intrinsic and extrinsic parameters, our approach encodes camera rays within a unified spatial framework, injecting spatially informed physical knowledge. For enhancing the diversity of generated instances, we leverage over 1,000 hours of driving scene data for model training. Regarding multi-condition control, we integrate 3D conditions—such as 3D bounding boxes and high-definition maps (HDmaps)—alongside image-wise textual descriptions. In conclusion, UniMLVG effectively supports text-to-video (T2V), image-to-video (I2V), and video-to-video (V2V) generation, demonstrating exceptional performance in both text-based and 3D condition-based editing, as illustrated in Fig. 1. The primary contributions:
@@
  • To enhance the motion coherence and the consistency across multiple views, we inject spatially physical knowledge through explicit perspective rays modeling.
@@ 2.2 Multi-view Video Generation
Compared to standard video generation, multi-view video generation remains relatively underexplored, due to the challenges in ensuring the consistency across perspectives and time series initial trials simply adopts the video diffusion architecture. For instance, MagicDrive [12] encodes high-level controls such as text and bounding boxes independently and flatten those conditions into an embedding sequence. In contrast, DreamForge [30] employs a ControlNet [62] to fuse multi-modality condition. Drive-WM [51] generates videos from a few views using a diffusion model in the first stage and then utilizes multi-view factorization to predict additional views conditioned on the pre-generated ones. DiVE [23] leverages the DiT structure within its diffusion model and applies a view-inflated attention mechanism to compute attention across features from all viewpoints. Extending the design of [50], DriveDreamer-2 [63] concatenates all images per frame into a large single image, and only cross-frame modules are adopted in the forward pass. Despite these approaches, they usually rely on a reference frame to generate subsequent content and the quality of the videos significantly degrades as the sequence length increases. Additionally, they struggle to address inconsistencies across views and timestamps. Based on the aforementioned issues, we introduce our unified model, which can provide precise control, generate high-quality videos and guarantee the spatial and temporal consistency.
@@ 3 Method
Figure 2: Overall framework of the model. left: The encoded reference frames are concatenated with the noisy latent as video latent and fed into $N$ UniMLVG blocks. The diverse conditions including image-level descriptions, camera pose, and 3D conditions are injected into each UniMLGV block and interact with the video latent to guide the generated contents. Finally, the model outputs the subsequent frames, which can then be used as the reference frame for the next autoregressive generation. Note that our model can produce driving video based on those conditions only, where the reference frames are not required. right: Details of the UniMLVG block. A UniMLVG block comprises three distinct sub-blocks to perform attention across different dimensions, while the different conditions are integrated into the video latent in different positions during the forward passing.
Fig. 2 illustrates the overall architecture of our proposed UniMLVG. Our method enhances the DiT-based image generation model [10] with two additional modules: the temporal module $\mathcal{T}$ and the cross-view module ${}\mathcal{C}$. UniMLVG is capable of generating extended street-view videos and performing text-based and 3D-conditioned editing by using a multi-task, multi-condition and multi-stage training strategy across diverse datasets. Importantly, we introduce explicit perspective modeling for the first time to incorporate physical spatial information, enabling smoother and more consistent multi-view autonomous driving scene videos.
@@ 3.1 Unified Framework
We choose the DiT-based model [10] as the backbone for image generation over alternatives like [38, 3, 11, 32] and extend the MM-DiT block, shown in the rigth of Figure 2 to our UniMLVG block to achieve an optimal balance of generation quality, scalability, model size, and flexibility in text-based control. For our diffusion model’s optimization, we employ the conditional rectified flows [26, 27, 10] loss as the primary objective:
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|v_{\theta}(z_{t},t,c)-(z_{0}-\epsilon)\right\|^{2}\right],$ (1)
where $v$ is the model, $z$ is the video latent, $\theta$ is the learnable parameters, $t$ is the timestep and $c$ is the condition.
@@ 3.2 Multi-task
We observe that when using only the video prediction task, the quality of the generated video deteriorates significantly after a few autoregressive iterations. We believe this is due to the model’s excessive reliance on reference frames, which leads to accumulated autoregressive errors. To address this issue, we propose a multi-task training strategy to improve long-term video quality and coherence. Specifically, we design four training tasks: video prediction (VP), image prediction (IP), video generation (VG), and image generation (IG). In VP, we utilize the embeddings of the first $k$ frames from $n$ viewpoints as reference frames, setting their timestep to 00. In IP, we mask 50% of the reference frames, requiring the model to use the remaining discrete reference information to generate images at the masked positions. In VG, the model generates the next $l$ frames of a multi-view video based solely on the given conditions, without any reference frames. In IG, we drop the temporal module to prevent the model from overly relying on temporal continuity, thus preserving its ability to maintain cross-view consistency. During training, the first three tasks are executed by sampling a mask $M\in\mathbb{R}^{T\times V}$ based on a predefined ratio, with the loss calculation excluding predictions of reference frames, as described below:
$\mathcal{L}=\mathbb{E}_{\epsilon\sim\mathcal{N}(0,I)}\left[\left\|(1-M)\odot(v_{\theta}(z_{t},t,c)-(z_{0}-\epsilon))\right\|^{2}\right].$ (2)
Notably, unlike DriveDreamer-2 [63], which considers only the VP and VG tasks, UniMLVG additionally incorporates the IP and IG tasks. As shown in Tab. 5 and Fig. 7, this contributes significantly to improving the quality of generated frames and mitigating autoregressive errors.
@@ 3.3 Multi-condition
To help the model grasp the physical dynamics of autonomous driving scenes, we introduce both local conditions (such as 3Dboxes and HDmaps) and global conditions (including view-specific text descriptions). Notably, we are the first to explicitly model camera parameters to incorporate physical spatial information for this task.
Local Conditions. We unify local conditions as image-based conditions. Similar to the approach used in T2I-Adapter [31], we introduce a lightweight image adapter for efficient processing. By projecting 3Dboxes or HDmaps and mapping the instance identity to the color space, we generate sparse images $I_{l}\in\mathbb{R}^{T\times V\times H^{\prime}\times W^{\prime}\times 3}$, matching the original image size. These conditional images are concatenated and input into the adapter to obtain multi-level features $C_{l}\in\mathbb{R}^{k\times T\times V\times H\times W\times C}$, which are then added to the latent representation at the corresponding layer.
@@
Figure 3: Field of view overlap between cameras over time.
Explicit Perspective Modeling. Apparently, there exist overlaps in the image contents between the adjacent views at the same moment and between the same view at different time. Moreover, as shown in Figure 3, there is also a significant field of view overlap between the current viewpoint of the moving vehicle and one of previous viewpoint. These overlaps help reduce uncertainty in generating scene images at new time steps. Meanwhile, modeling camera poses explicitly has proven effective in recent works [13, 41] that generate multi-view images of static scenes. We believe this is also applicable for generating consistent dynamic driving scenes. Specifically, we set the optical center of the forward-facing camera in the first frame as the origin to establish a unified coordinate system for all viewpoints across different time steps. We encode the camera ray representation [40] to produce ray features, which are added as spatial information to the video latent. The formula for ray coordinate encoding is as follows:
$\mathcal{P}(c)=$ $[\sin(2^{0}\pi c),\cos(2^{0}\pi c),\sin(2^{1}\pi c),$
$\cos(2^{1}\pi c),\dots,\sin(2^{j-1}\pi c),\cos(2^{j-1}\pi c)]$
$Enc(\mathbf{v})=$ $MLP(Concat(\mathcal{P}(\mathbf{v}_{0}),\mathcal{P}(\mathbf{v}_{1}),\dots,\mathcal{P}(\mathbf{v}_{k}))),$
where $\mathcal{P}$ denotes the encoding function applied to each value of $\mathbf{v}$ (the center and direction of ray), $j$ represents the encoding dimension (set to $8$ in this paper).
@@
Ours $\surd$ $\surd$ 20s 5.8 36.1 22.50 70.81 29.12
@@ 3.4 Multi-stage
UniMLVG utilizes a wide range of currently available datasets, including both single-view and multi-view street scene videos, to impart driving scene priors to the model. Considering that single-view and multi-view data differ in the number of viewpoints and annotation information, we implement a multi-stage training strategy to ensure stable and efficient model convergence. It is worth noting that our multi-stage training does not follow the typical approach of first training on images and then fine-tuning on videos as in previous works. Instead, all stages of our training are conducted using video data.
@@
Stage II. Infusing the model with the ability to generate from multiple viewpoints and effectively follow conditional inputs. We train $\mathcal{C}$ and $\mathcal{T}$ using several multi-view datasets [5, 44, 53]. These datasets provide multi-view videos, camera calibration parameters, and 3D annotations. During this phase, only the temporal and cross-view modules are trained, while the backbone remains frozen.
Stage III. To further improve generation quality, we perform full fine-tuning in this phase. With the model’s generation capabilities well-developed from the first two stages, this phase typically requires 1 or 2 training epochs.
@@ 4.1 Experiment Details
Datasets. We use the single-view dataset OpenDV-Youtube [58] and the multi-view datasets nuScenes [5], Waymo [44], and Argoverse2 [53]. OpenDV-Youtube is exclusively used for Stage I, while nuScenes, Waymo, and Argoverse2 are divided into training and validation sets following their original splits. The training duration totals $1,498$ hours, comprising $1,486$ hours from OpenDV-Youtube, $4.6$ hours from nuScenes, $4.4$ hours from Waymo, and $3.1$ hours from Argoverse2. We leverage available dataset annotations, including 3Dboxes, HDmaps and camera parameters. Meanwhile, nuScenes with $12$ Hz interpolated annotations [2] is used. Additionally, text descriptions for all frames and views are generated at $2$ Hz using [49].
Evaluation Metrics. To assess the effectiveness of our method in terms of realism, continuity, and precise control, we selected four key metrics to compare against existing multi-view image and video generation methods. For realism, we use the widely recognized Fréchet Inception Distance (FID) [17]. To estimate temporal coherence in our videos, we measure consistency using Fréchet Video Distance (FVD) [46]. For a fair comparison, we conduct evaluation following [12], using $150$ scenes from the nuScenes validation set, with $6$ viewpoints per scene and $16$ frames per viewpoint, totaling $900$ videos for FVD and FID computation. For controllability, we evaluate two perception tasks: 3D object detection [28] and BEV segmentation [65], following the approach MagicDrive [12].
@@ 4.2 Experiment Results
Quantitative Results. We report quantitative experimental metrics on the nuScenes validation set, as shown in Tab. 1. Overall, our model achieves quite promising results, with a significant improvement of 48.2% in FID and 35.2% in FVD compared to the second-best method. As demonstrated in the ablation study 4.3, the multi-task, multi-stage training strategies and explicit perspective modeling make significant contributions. In terms of temporal consistency, we achieve a 61.8% improvement over DiVE [23], which also uses the DiT [18] architecture. Compare to DriveDreamer-2 [63], UniMLVG achieve a 35.2% improvement in FVD. Moreover, DriveDreamer-2 can only produce 7s video while UniMLVG can generate longer video up to $20$s, almost three times as longer as DriveDreamer-2. Additionally, unlike DriveDreamer-2 [63], UniMLVG does not stitch multiple views into a single image for generation, which reduces memory usage and training time. On contrast, our cross-view module can easily handle the different numbers of viewpoints presented in different datasets using view mask. In terms of condition adherence, UniMLVG achieve the SOTA results, improving $\text{mIoU}_{road}$ and $\text{mIoU}_{vehicle}$ by 8.8% and 7.8%, respectively, over the second-best method. For mAP, we outperform all other methods except DiVE [23]. In contrast to DiVE, UniMLVG does not assign distinct classifier-free guidance scales for each condition, significantly reducing inference time. Furthermore, rather than using the parameter-heavy ControlNet [62], UniMLVG employs a lightweight adapter, enabling more flexible and efficient condition integration.
@@
$\surd$ 49.30 8.80
$\surd$ 49.18 8.86
$\surd$ $\surd$ 50.13 8.91
$\surd$ $\surd$ $\surd$ 50.76 8.74
$\surd$ $\surd$ $\surd$ 44.43 6.78
@@
Controllability. Our model supports 3D condition control as well as text-based control capabilities. In Figure 1(c), we generate realistic scenes using 3D conditions obtained from the simulation engine CALAR [9]. Notably, different from style transfer methods, our generation is not strictly constrained to transforming the content of simulated scenes. Instead, our model takes the general 3D information such as 3D bounding boxes and HDMaps as inputs and leverages real-world knowledge to generate plausible scenes that align with the distribution of the training data. Specifically, elements such as roads, mountains, and trees are generated with appearances that more accurately reflect how they would look in the real world. For textual control, the model illustrates strong cross-distribution transfer and impressive editing capabilities under extreme transformations. As shown in Figures 1(d) and 4, even though the nuScenes dataset does not contain any snowy, our model can still edit the weather of nuScene multi-view videos to snowy in both daytime and nightime. Additionally, the generated videos accurately present real-world details such as snow accumulating on the roadside, bare trees during winter, and reflective road surfaces due to rain.